[Brahms-l] Shift report 20040131 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sat Jan 31 2004 - 15:57:08 EST
------------------------------------------------------------
Date: Sat Jan 31 15:57:07 2004, Shift: 08:00-16:00
Supervisor: sl, with: 
Log book#15 pages 46 - 46

Shift Summary:
------------------------------------------------------------
1. Cryo problem in the morning, one long store 9am - 2.30pm new beam ready for data around 3.30
2.Finished ~500k tr2 events at 6deg 1723B (9978-9979)
  did      ~400k tr2 events at 6deg 1219B (9980-9982) 
  and      ~700k tr3 events at 40deg 1050A(9978-9982)
3. No detector problems.
            
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 31 15:58:03 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 31 2004 - 15:58:19 EST